Fermi National Laboratory


Accelerator Update for July 15, 2005 - July 18, 2005

Friday July 15
The day shift began with Operations monitoring store 4276, the Recycler Ring (RR) stash at 63.9 E10, the Antiproton Source (Pbar) stack at 11.7 E10, and with MiniBooNE, NuMI, and the Meson MCenter taking beam.

Store 4276 was terminated at 10:00 PM.

D0 went into access at 10:12 PM for approximately 90 minutes.

Saturday July 16
The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE , NuMI, and MCenter taking beam.

Shot setup begins at 12:21 AM.

Operations established store 4278 at 2:37 AM with an initial luminosity of 90.39E30. Pbar resumed stacking, beam returns to MCenter.

Operations put the Recycler into shot setup at 6:19 AM to stash antiprotons.

Pbar returns to stacking at 7:21 AM.

Operations put the Recycler into shot setup at 12:28 PM to stash antiprotons.

Recycler shots were completed at 2:06 PM, Operations returned to stacking, they also returned beam to MCenter.

At 2:27 PM, problems with Main Injector RF station 16 (MIRF16) will require a tunnel access to repair, the station has been removed from the loop until a tunnel access can be made for repair.

Sunday July 17
The midnight shift began with Operations monitoring store 4278, the Recycler Ring (RR) stash at 98.4 E10, the Pbar stack at 53.2 E10, and with MiniBooNE, NuMI, and MCenter taking beam.

Store 4278 was terminated at 5:03 AM.

Shot setup begins at 5:15 AM.

Operations established store 4279 at 7:45 AM with an initial luminosity of 100.07E30. Pbar resumed stacking, beam returns to MCenter.

Operations put the Recycler into shot setup at 12:23 PM to stash antiprotons.

Recycler shots were completed at 1:34 PM. Pbar has returned to stacking, Mcenter is once again taking beam.

Operations once again puts the Recycler into shot setup at 7:43 PM to stash antiprotons.

Recycler shots were completed at 8:52 PM.

Operations returned to stacking.

At 8:47 PM a MiniBoone quad power supply (E:Q874) began tripping its wall breaker, fans were taken out to the service building (MI-12) to try and alleviate the trips.

Monday July 18
The midnight shift began with Operations monitoring store 4279, monitoring the stash, Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

E:Q874 continues to trip during the midnight shift, experts will look at it in the morning.

The Plan for Monday
Is to terminate the store at 11:00 AM and then go into shot setup.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/18/2005   email Fermilab